For a reaction with ΔH∘=−25 kcal/mol and ΔS∘=0.05 kcalmol−1K−1, calculate the equilibrium constant at: (1.) 30 ∘C and (2.) 150 ∘C.

Express your answers using two significant figures separated by a comma.Homework for Unit #4: Chapter 5 Problem 5.22 80115 Review | Constants Periodic Table Part A For a reaction with AH = -25 kca
Answer #1

AH = -25 Keal I mol 45°3005Kcas/not र T = 300- 30+र+3 = 303K = 1500-150+01343K AGAR - TA5 = -25 kcal/mole - 303 XO 05 45 -40